Solve for x: logbase2(x-6)+logbase2(x)=4
Answer by edjones(8007)

logbase2(x-6)+logbase2(x)=4
2^log[2](x-6)(x)=2^4
x^2-6x=16
x^2-6x-16=0
(x-8)(x+2)=0
x=8
x cannot be -2 because the domain of logs dont include negative numbers.
.
Ed